OpenStudy (anonymous):

Can someone please teach me how to multiply mixed numbers? 1 1/4*4 4/5

OpenStudy (nikato):

you can start by changing them to improper fractions. do you know how?

OpenStudy (anonymous):

yeah

OpenStudy (nikato):

ok. can you change both fractions into improper?

OpenStudy (anonymous):

1 1/4 =5/4 and 4 4/5 =24/5

OpenStudy (anonymous):

then

OpenStudy (nikato):

ok. now multiply them